OpenStudy (anonymous):

tom picked 10 beads from the bag and recorded his results below: red:2 blue:2 yellow:5 purple:1 if the bag contained 75 beads, how many of the following beads should you pick from the bag? (use the theoretical probability)

OpenStudy (anonymous):

2/10 beads were red so 1/5 are red 1/5 * 75 = 15 reds 2/10 beads were blue so 1/5 are blue 1/5*75 = 15 blues 5/10 beads were yellow so 1/2 are yellow 1/2 * 75 = 37.5 so either 37 or 38 yellow 1/10 beads were purple so 1/10 are purple 1/10 * 75 = 7.5 so either 7 or 8 purple

OpenStudy (anonymous):

thank you soososososo much(: